A value of Ksp for Ca(OH)2 from one source is 5.50 x 10-5 .

a. Calculate the theoretical molar solubility of Ca(OH)2.

b. Calculate the expected pH of a saturated (and filtered) Ca(OH)2 solution.
Answer #1

A)

At equilibrium:

Ca(OH)2 <----> Ca2+ + 2 OH-

   s 2s

Ksp = [Ca2+][OH-]^2

5.5*10^-5=(s)*(2s)^2

5.5*10^-5= 4(s)^3

s = 2.396*10^-2 M

Answer: 2.40*10^-2 M

B)

[OH-] = 2s

= 2*2.396*10^-2 M

= 4.792*10^-2 M

use:

pOH = -log [OH-]

= -log (4.792*10^-2)

= 1.3195

use:

PH = 14 - pOH

= 14 - 1.3195

= 12.6805

Answer: 12.68